element A has two isotopes. The first isotope is pressnt 18.48% of the time and has a mass of 135.24. The second isotope has a mass of 116.1. calculate the atomic mass of element A
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. Thanks in advance.
Answer #2

To calculate the atomic mass of element A, we need to consider the masses and relative abundances of its isotopes.

Let's assume the first isotope has a mass of x and is present 18.48% of the time. The second isotope has a mass of 116.1 and is present (100 - 18.48)% = 81.52% of the time.

We can set up the following equation based on the weighted average:

(x * 18.48%) + (116.1 * 81.52%) = Atomic mass of element A

Simplifying the equation:

0.1848x + 95.02672 = Atomic mass of element A

To solve for x, we subtract 95.02672 from both sides:

0.1848x = Atomic mass of element A - 95.02672

Now, we substitute the mass value of the first isotope (135.24) for x:

0.1848(135.24) = Atomic mass of element A - 95.02672

24.99152 = Atomic mass of element A - 95.02672

To find the atomic mass of element A, we add 95.02672 to both sides:

Atomic mass of element A = 24.99152 + 95.02672

Atomic mass of element A = 120.01824

Therefore, the atomic mass of element A is approximately 120.018 (to 3 decimal places).


answered by: mervetokaz